# Markov spin models for image generation : explicit large deviations with respect to the number of pixels

### v1

@@ III.2.1 Generating function ${\cal G}^{[\vec{\sigma}]}_{t}(h,g)$ for any finite $N$
It is thus convenient to introduce the notation
@@
III.2.2 Large deviations for large $N$ : scaled cumulant generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$
For large $N$, the generating function ${\cal G}_{t}^{[M_{0}]}(h,g)$ of Eq. 57 is governed by the exponential behavior in $N$
${\cal G}_{t}^{[M_{0}]}(h,g)\mathop{\propto}_{N\to+\infty}e^{N{\cal F}_{t}^{[m_{0}]}(h,g)}$ (58)
where the scaled cumulant generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$ reads in terms of the intensive magnetization $m_{0}=\frac{M_{0}}{N}$ of Eq. 34
@@ III.2.2 Large deviations for large $N$ : scaled cumulant generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$
In particular, the values of the first-order derivatives for $h=0=g$ reproduce the typical values $(\hat{m},\hat{q})$ of Eq. 47 as it should
$\partial_{h}{\cal F}_{t}^{[m_{0}]}(h,g)\bigg{|}_{h=0=g}$ $=m_{0}\tanh\mu_{t}=m_{0}\lambda^{t}=\hat{m}$
$\partial_{g}{\cal F}_{t}^{[m_{0}]}(h,g)\bigg{|}_{h=0=g}$ $=\tanh\mu_{t}=\lambda^{t}=\hat{q}$ (64)
III.2.3 Legendre transformation between the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ and the scaled cumulant generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$
A central result in the theory of large deviations is that the rate function and the scaled cumulant generating function contain the same information, as we now recall for the present setting.
@@
yields via the saddle-point evaluation for large $N\to+\infty$ of the integrals over $m$ and $q$ that the scaled cumulant generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$ is related to the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ via the double Legendre transform that involves the first derivatives of Eqs 41 and 42
@@
while the reciprocal Legendre transformation involves the first derivatives of Eqs 60 and 60
${\cal I}_{t}^{[m_{0}]}(m,q)$ $=hm+gq-{\cal F}_{t}^{[m_{0}]}(h,g)$
@@
III.3 Large deviations properties of the intensive overlap $q$ alone
III.3.1 Explicit rate function $i_{t}(q)$ for the intensive overlap $q$ alone
$i_{t}^{[m_{0}]}(q)=\min_{m}\bigg{(}{\cal I}_{t}^{[m_{0}]}(m,q)\bigg{)}$ (68)
The derivative of Eq. 41
leads to the optimal magnetization
$m^{opt}=m_{0}q$ (70)
that can be plugged into the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ of Eq. 50 to obtain the rate function
$i_{t}^{[m_{0}]}(q)$ $={\cal I}_{t}^{[m_{0}]}(m^{opt}=m_{0}q,q)$ (71)
that is independent of the initial magnetization $m_{0}$ as expected.
At the typical value ${\hat{q}}=\lambda^{t}$, the rate function $i_{t}(q)$ vanishes $i_{t}({\hat{q}})=0$ together with its derivative
III.3.2 Scaled cumulant generating function $f_{t}(g)$ for the overlap $q$ alone
For the single variable $q$, the Legendre transformation between the rate function $i_{t}(q)$ and the scaled cumulant generating function $f_{t}(g)$ analogous to Eq. 66 involves the derivative of Eq. 72
$f_{t}(g)$ $=gq-i_{t}(q)$
with the reciprocal Legendre transformation
$i_{t}(q)$ $=gq-f_{t}(g)$
$q$ $=\frac{df_{t}(g)}{dg}=\tanh(g+\mu_{t})$ (75)
@@ III.4.1 Rate function $I_{t}^{[m_{0}]}(m)$ for the intensive magnetization $m$ alone
The rate function $I_{t}^{[m_{0}]}(m)$ for the magnetization $m$ alone corresponds to the optimization of the joint rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ over the overlap $q$
$I_{t}^{[m_{0}]}(m)=\min_{q}\bigg{(}{\cal I}_{t}^{[m_{0}]}(m,q)\bigg{)}$ (76)
The derivative of Eq. 42
leads to the second-order equation in $q$
$=4\lambda^{t}(q^{2}+1-m^{2}-m_{0}^{2})-4(1+\lambda^{2t})(q-mm_{0})$
that becomes after factoring out the global factor $(4\lambda^{t})$
00 0 $=q^{2}+1-m^{2}-m_{0}^{2}-(\lambda^{-t}+\lambda^{t})(q-mm_{0})$ (79)
The discriminant
leads to the two solutions
$q_{\pm}=\frac{\lambda^{t}+\lambda^{-t}\pm\sqrt{\Delta}}{2}$ (81)
Since their sum is bigger than 2
the solution $q_{+}$ is bigger than unity $q_{+}>1$ and is thus outside the interval $[-1,+1]$ of physical overlaps. The physical solution is thus $q_{-}$, and one needs to evaluate the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ of Eq. 50 for this special value $q=q_{-}$ using Eq. 41
that can be used to rewrite the ratio
so that the first contribution of Eq. 83 can be rewritten as
and leads to the rate function of Eq. 83 with Eq. 86
Since $q={\mathring{q}}_{t}^{[m_{0}]}(m)$ solves the optimization problem of Eq. 77, the derivative of the rate function $I_{t}^{[m_{0}]}(m)$ of Eq. 88 with respect to $m$ only involves the partial derivative $\partial_{m}{\cal I}_{t}^{[m_{0}]}(m,q)$ of Eq. 41
$\partial_{m}I_{t}^{[m_{0}]}(m)$ $=\partial_{m}{\cal I}_{t}^{[m_{0}]}(m,q)\bigg{|}_{q={\mathring{q}}_{t}^{[m_{0}]}(m)}+\left[\partial_{m}{\mathring{q}}_{t}^{[m_{0}]}(m)\right]\partial_{q}{\cal I}_{t}^{[m_{0}]}(m,q)\bigg{|}_{q={\mathring{q}}_{t}^{[m_{0}]}(m)}=\partial_{m}{\cal I}_{t}^{[m_{0}]}(m,q)\bigg{|}_{q={\mathring{q}}_{t}^{[m_{0}]}(m)}+0$ (89)
that can be rewritten using Eq. 86
When $m$ takes its typical value ${\hat{m}}=\lambda^{t}m_{0}$ of Eq. 47, the optimal value ${\mathring{q}}_{t}^{[m_{0}]}(m={\hat{m}})$ Eq. 87 coincides with the typical value $\hat{q}=\lambda^{t}$
with the corresponding vanishing of the rate function of Eq. 88 and of its derivative of Eq. 90
$I_{t}^{[m_{0}]}({\hat{m}})$ $=0$
$\partial_{m}I_{t}^{[m_{0}]}\bigg{|}_{m={\hat{m}}}$ $=0$ (92)
@@ III.5 Discussion
In summary, the large deviations properties of the intensive magnetization $m=\frac{M}{N}$ and of the intensive overlap $q=\frac{Q}{N}$ produced by the forward propagator associated to the initial condition $\vec{\sigma}$ of intensive magnetization $m_{0}$ have been analyzed in detail via explicit rate functions and via explicit scaled cumulant generating functions that are related by Legendre transformations. In the three following sections, we analyze the consequences for the backward-generative dynamics for various initial conditions $P^{ini}(.)$
@@ IV Analysis of the backward dynamics for a single initial condition $\vec{\sigma}$
In this section, we consider that the initial condition $P^{ini}(\vec{S})$ reduces the single image $\vec{\sigma}$
@@
IV.1 Direct ink with the Doob conditioning of the forward Markov process with a finite horizon $T$
@@ IV.2 Explicit backward-generated distribution $B^{Gene}_{t}(\vec{S}(t))$ for any finite $N$
In order to compute Eq. 97 and to compare with the forward propagator $\langle\vec{S}(t)|{\cal W}^{t}|\vec{\sigma}\rangle$, let us focus on the ratio
@@
Since the forward propagator $\langle\vec{S}(t)|{\cal W}^{t}|\vec{\sigma}\rangle$ also depends only on $Q(\vec{S}(t);\vec{\sigma})$ via Eq. 18, the final result is that the backward-generated distribution $B^{Gene}_{t}(\vec{S}(t))$ reduces to
$B^{Gene}_{t}(\vec{S}(t))=\left(\frac{(1+\lambda^{t})(1-\lambda^{2T-t})}{2(1-\lambda^{2T})}\right)^{\frac{N+Q(\vec{S};\vec{\sigma})}{2}}\left(\frac{(1-\lambda^{t})(1+\lambda^{2T-t})}{2(1-\lambda^{2T})}\right)^{\frac{N-Q(\vec{S};\vec{\sigma})}{2}}$ (108)
@@
$b_{t}^{Gene}(Q)$ $=\sum_{\vec{S}(t)}\delta_{Q,\sum_{n=1}^{N}S_{n}(t)\sigma_{n}}B^{Gene}_{t}(\vec{S}(t))$ (109)
$=\left(\frac{(1+\lambda^{t})(1-\lambda^{2T-t})}{2(1-\lambda^{2T})}\right)^{\frac{N+Q(\vec{S};\vec{\sigma})}{2}}\left(\frac{(1-\lambda^{t})(1+\lambda^{2T-t})}{2(1-\lambda^{2T})}\right)^{\frac{N-Q(\vec{S};\vec{\sigma})}{2}}\sum_{\vec{S}(t)}\delta_{Q,\sum_{n=1}^{N}S_{n}(t)\sigma_{n}}$
$=\left(\frac{(1+\lambda^{t})(1-\lambda^{2T-t})}{2(1-\lambda^{2T})}\right)^{\frac{N+Q(\vec{S};\vec{\sigma})}{2}}\left(\frac{(1-\lambda^{t})(1+\lambda^{2T-t})}{2(1-\lambda^{2T})}\right)^{\frac{N-Q(\vec{S};\vec{\sigma})}{2}}\frac{N!}{\frac{N+Q}{2}!\frac{N-Q}{2}!}$

### v2

@@ III.3.1 Generating function ${\cal G}^{[\vec{\sigma}]}_{t}(h,g)$ for any finite $N$
It is then convenient to introduce the notation
@@
The series expansion of ${\cal G}^{[\vec{\sigma}]}_{t}(h,g)$ of Eq. 55 with respect to the two variables $(h,g)$
${\cal G}^{[M_{0}]}_{t}(h,g)$ $\equiv\sum_{M}\sum_{Q}e^{hM+gQ}P^{[\vec{\sigma}]}_{t}(M,Q)$ (61)
$=1+\sum_{M}\sum_{Q}\left[hM+gQ+\frac{h^{2}M^{2}+2hgMQ+g^{2}Q^{2}}{2}+...\right]P^{[\vec{\sigma}]}_{t}(M,Q)$
$\equiv 1+h{\cal M}_{1}(t)+g{\cal Q}_{1}(t)+\frac{h^{2}}{2}{\cal M}_{2}(t)+\frac{g^{2}}{2}{\cal Q}_{2}(t)+hg{\cal C}_{1,1}(t)+...$
involves the various moments and correlations of the joint distribution $P^{[\vec{\sigma}]}_{t}(M,Q)$
${\cal M}_{k}(t)$ $\equiv\sum_{M}\sum_{Q}M^{k}P^{[\vec{\sigma}]}_{t}(M,Q)$
${\cal Q}_{l}(t)$ $\equiv\sum_{M}\sum_{Q}Q^{l}P^{[\vec{\sigma}]}_{t}(M,Q)$
${\cal C}_{k,l}(t)$ $\equiv\sum_{M}\sum_{Q}M^{k}Q^{l}P^{[\vec{\sigma}]}_{t}(M,Q)$ (62)
while the generating function of the various cumulants and connected correlations appear in the series expansion of its logarithm
$\ln\left({\cal G}^{[\vec{\sigma}]}_{t}(h,g)\right)$ $=\ln\left(1+h{\cal M}_{1}(t)+g{\cal Q}_{1}(t)+\frac{h^{2}}{2}{\cal M}_{2}(t)+\frac{g^{2}}{2}{\cal Q}_{2}(t)+hg{\cal C}_{1,1}(t)+...\right)$ (63)
$=h{\cal M}_{1}(t)+g{\cal Q}_{1}(t)+\frac{h^{2}}{2}\left({\cal M}_{2}(t)-{\cal M}_{1}^{2}(t)\right)+\frac{g^{2}}{2}\left({\cal Q}_{2}(t)-{\cal Q}_{1}^{2}(t)\right)+hg\left({\cal C}_{1,1}(t)-{\cal M}_{1}{\cal Q}_{1}\right)+...$
III.3.2 Large deviations for large $N$ : generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$ of scaled cumulants
For large $N$, the logarithm of the generating function ${\cal G}_{t}^{[M_{0}=Nm_{0}]}(h,g)$ of Eq. 60 is dominated by the extensive behavior with respect to the size $N$
$\ln\left({\cal G}^{[M_{0}=Nm_{0}]}_{t}(h,g)\right)\mathop{\propto}_{N\to+\infty}N{\cal F}_{t}^{[m_{0}]}(h,g)$ (64)
where the generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$ of the scaled cumulants
${\cal F}_{t}^{[m_{0}]}(h,g)$ $=\lim_{N\to+\infty}\left[\frac{\ln\left({\cal G}^{[\vec{\sigma}]}_{t}(h,g)\right)}{N}\right]$
$=\lim_{N\to+\infty}\left[h\frac{{\cal M}_{1}(t)}{N}+g\frac{{\cal Q}_{1}(t)}{N}+\frac{h^{2}}{2}\left(\frac{{\cal M}_{2}(t)-{\cal M}_{1}^{2}(t)}{N}\right)+\frac{g^{2}}{2}\left(\frac{{\cal Q}_{2}(t)-{\cal Q}_{1}^{2}(t)}{N}\right)+hg\left(\frac{{\cal C}_{1,1}(t)-{\cal M}_{1}{\cal Q}_{1}}{N}\right)+...\right]$
reads in terms of the intensive magnetization $m_{0}=\frac{M_{0}}{N}$
@@
In particular for $h=0=g$, the first-order derivatives correspond to the averaged values of the intensive magnetization and of the intensive overlap that appear at first-order in the series expansion of Eq. III.3.2
$\partial_{h}{\cal F}_{t}^{[m_{0}]}(h,g)\bigg{|}_{h=0=g}$ $=m_{0}\tanh\mu_{t}=m_{0}\lambda^{t}=\lim_{N\to+\infty}\left(\frac{{\cal M}_{1}(t)}{N}\right)=\hat{m}$
$\partial_{g}{\cal F}_{t}^{[m_{0}]}(h,g)\bigg{|}_{h=0=g}$ $=\tanh\mu_{t}=\lambda^{t}=\lim_{N\to+\infty}\left(\frac{{\cal Q}_{1}(t)}{N}\right)=\hat{q}$ (71)
that are in agreement with the typical values $(\hat{m}=\lambda^{t}m_{0},\hat{q}=\lambda^{t})$ of Eq. 49 as it should, while the second-order derivatives lead to the rescaled variances of the magnetization and of the overlap, and to their rescaled connected correlation that appear at second-order in the series expansion of Eq. III.3.2
$\partial_{h}^{2}{\cal F}_{t}^{[m_{0}]}(h,g)\bigg{|}_{h=0=g}$ $=\frac{1}{\cosh^{2}\left[\mu_{t}\right]}=1-\lambda^{2t}=\lim_{N\to+\infty}\left(\frac{{\cal M}_{2}(t)-{\cal M}_{1}^{2}(t)}{N}\right)$
$\partial_{g}^{2}{\cal F}_{t}^{[m_{0}]}(h,g)\bigg{|}_{h=0=g}$ $=\frac{1}{\cosh^{2}\left[\mu_{t}\right]}=1-\lambda^{2t}=\lim_{N\to+\infty}\left(\frac{{\cal Q}_{2}(t)-{\cal Q}_{1}^{2}(t)}{N}\right)$
$\partial_{g}\partial_{h}{\cal F}_{t}^{[m_{0}]}(h,g)\bigg{|}_{h=0=g}$ $=\frac{m_{0}}{\cosh^{2}\left[\mu_{t}\right]}=m_{0}\left(1-\lambda^{2t}\right)=\lim_{N\to+\infty}\left(\frac{{\cal C}_{1,1}(t)-{\cal M}_{1}{\cal Q}_{1}}{N}\right)$ (72)
These values are in agreement with their alternative evaluations from the joint Gaussian approximation of Eq. 54 for the pair $(m,q)$ around the typical value $(\hat{m}=\lambda^{t}m_{0},\hat{q}=\lambda^{t})$.
III.3.3 Legendre transformation between the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ and the scaled cumulant generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$
A central result in the theory of large deviations is that the rate function and the scaled cumulant generating function are directly related via Legendre transformations, as we now recall for the present setting involving two variables.
@@
yields, via the saddle-point evaluation for large $N\to+\infty$ of the integrals over the two variables $m$ and $q$, that the scaled cumulant generating function ${\cal F}_{t}^{[m_{0}]}(h,g)$ is related to the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ via the double Legendre transform that involves the first derivatives of Eqs 44 and 45
@@
while the reciprocal Legendre transformation involves the first derivatives of Eqs 67 and 68
${\cal I}_{t}^{[m_{0}]}(m,q)$ $=hm+gq-{\cal F}_{t}^{[m_{0}]}(h,g)$ (75)
@@ III.4.1 Rate function $I_{t}^{[m_{0}]}(m)$ for the intensive magnetization $m$ alone
The rate function $I_{t}^{[m_{0}]}(m)$ for the magnetization $m$ alone corresponds to the optimization of the joint rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ over the overlap $q\in[-1,+1]$
$I_{t}^{[m_{0}]}(m)=\min_{q\in[-1,+1]}\bigg{(}{\cal I}_{t}^{[m_{0}]}(m,q)\bigg{)}$ (76)
whose properties are described in Appendix B, while the Gaussian typical fluctuations of order $\frac{1}{\sqrt{N}}$ of the magnetization $m$ around its typical value ${\hat{m}}=\lambda^{t}m_{0}$ can be directly obtained from the integration over $q$ of the joint Gaussian approximation of Eq. 54
[…]
in agreement with the rescaled variance of the magnetization of Eq. 72.
@@ III.5 Discussion
In summary, the joint statistical properties of the overlap $Q$ and of the magnetization $M$ produced by the forward propagator associated to any given initial condition $\vec{\sigma}$ of magnetization $M_{0}$ have been characterized for any time $t$, both for any finite number $N$ of pixels and in the limit $N\to+\infty$, where the large deviations are governed by explicit rate functions and explicit scaled cumulant generating functions. In the three following sections, the goal will be to analyze the consequences for the backward-generative dynamics for various initial conditions $P^{ini}$.
@@ IV Analysis of the backward dynamics for a single initial condition $\vec{\sigma}$
In this section, we consider the case where the initial condition $P^{ini}(\vec{S})$ reduces the single image $\vec{\sigma}$
@@
IV.1 Direct link with the Doob conditioning of the forward Markov process with a finite horizon $T$
@@ IV.2 Explicit backward-generated distribution $B^{Gene}_{t}(\vec{S}(t))$ for any finite $N$
In order to compute the backward generated distribution $B^{Gene}_{t}(\vec{S}(t))$ of Eq. 82 and to compare with the forward propagator $\langle\vec{S}(t)|{\cal W}^{t}|\vec{\sigma}\rangle$, let us focus on the ratio
@@
Since the forward propagator $\langle\vec{S}(t)|{\cal W}^{t}|\vec{\sigma}\rangle$ also depends only on the overlap $Q(\vec{S}(t);\vec{\sigma})$ via Eq. 20
$\langle\vec{S}(t)|{\cal W}^{t}|\vec{\sigma}\rangle=\left(\frac{1+\lambda^{t}}{2}\right)^{\frac{N+Q(\vec{S}(t);\vec{\sigma})}{2}}\left(\frac{1-\lambda^{t}}{2}\right)^{\frac{N-Q(\vec{S}(t);\vec{\sigma})}{2}}$ (93)
the final result is that the backward-generated distribution $B^{Gene}_{t}(\vec{S}(t))$
$B^{Gene}_{t}(\vec{S}(t))$ $=\left(\frac{(1+\lambda^{t})(1-\lambda^{2T-t})}{2(1-\lambda^{2T})}\right)^{\frac{N+Q(\vec{S}(t);\vec{\sigma})}{2}}\left(\frac{(1-\lambda^{t})(1+\lambda^{2T-t})}{2(1-\lambda^{2T})}\right)^{\frac{N-Q(\vec{S}(t);\vec{\sigma})}{2}}$ (94)
$=\left(\frac{1+\lambda^{t}\left[\frac{1-\lambda^{2(T-t)}}{1-\lambda^{2T}}\right]}{2}\right)^{\frac{N+Q(\vec{S}(t);\vec{\sigma})}{2}}\left(\frac{1-\lambda^{t}\left[\frac{1-\lambda^{2(T-t)}}{1-\lambda^{2T}}\right]}{2}\right)^{\frac{N-Q(\vec{S}(t);\vec{\sigma})}{2}}$
has exactly the same form as the forward propagator of Eq. 93 except for the replacement
$\lambda^{t}\to\lambda^{t}\left[\frac{1-\lambda^{2(T-t)}}{1-\lambda^{2T}}\right]$ (95)
@@
$b_{t}^{Gene}(Q)$ $\equiv\sum_{\vec{S}(t)}\delta_{Q,\sum_{n=1}^{N}S_{n}(t)\sigma_{n}}B^{Gene}_{t}(\vec{S}(t))$ (96)
$=\left(\frac{1+\lambda^{t}\left[\frac{1-\lambda^{2(T-t)}}{1-\lambda^{2T}}\right]}{2}\right)^{\frac{N+Q(\vec{S}(t);\vec{\sigma})}{2}}\left(\frac{1-\lambda^{t}\left[\frac{1-\lambda^{2(T-t)}}{1-\lambda^{2T}}\right]}{2}\right)^{\frac{N-Q(\vec{S}(t);\vec{\sigma})}{2}}\sum_{\vec{S}(t)}\delta_{Q,\sum_{n=1}^{N}S_{n}(t)\sigma_{n}}$
$=\left(\frac{1+\lambda^{t}\left[\frac{1-\lambda^{2(T-t)}}{1-\lambda^{2T}}\right]}{2}\right)^{\frac{N+Q(\vec{S}(t);\vec{\sigma})}{2}}\left(\frac{1-\lambda^{t}\left[\frac{1-\lambda^{2(T-t)}}{1-\lambda^{2T}}\right]}{2}\right)^{\frac{N-Q(\vec{S}(t);\vec{\sigma})}{2}}\frac{N!}{\frac{N+Q}{2}!\frac{N-Q}{2}!}$
that has the same form as the forward result of Eq. 27 up to the replacement of Eq. 95.